How many liters of a 10% acid solution must be mixed with a 50% acid solution in order to obtain 25 litres of a 30% solution?
Answer by richwmiller(17219)

x+y=25
.1*x+.5y=.3*25
x=12.5 and y=12.5
.1*12.5+.5*12.5=.3*25=7.5
check
1.25+6.25=7.5
ok